OpenStudy (anonymous):

3. What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10?

OpenStudy (anonymous):

Can you answer my question? This is not a public dating site. -_-

jimthompson5910 (jim_thompson5910):

You need the first term of the sequence an = a1 + d(n-1) 13 = a1 - 10(12-1) 13 = a1 - 10(11) 13 = a1 - 110 13+110 = a1 123 = a1 a1 = 123 So the first term is 123

jimthompson5910 (jim_thompson5910):

With the first term, the last term and the number of terms, you can figure out what the first 12 terms add to

jimthompson5910 (jim_thompson5910):

using this formula S = (n*(a1 + an))/2 S = (12*(123 + 13))/2 ... plug in n = 12, a1 = 123 and an = 13 S = ??

jimthompson5910 (jim_thompson5910):

karinewoods17, just ignore POOPSCOOPER153 since he looks like a troll to me

OpenStudy (anonymous):

Thank you so much! I didn't know how to find the first term. You are so helpfull. I wish I could give you 1,000,000,000,000... medals:)

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

no not really, she's just grateful for me actually helping (instead of trolling)

OpenStudy (anonymous):

Exactly! Ugh. stupid immature boys.

OpenStudy (anonymous):

Is the answer 816?

jimthompson5910 (jim_thompson5910):

you got it, the sum of the first 12 terms is 816

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw